# VD3D: Taming Large Video Diffusion Transformers for 3D Camera Control

### v2

@@ Abstract
Modern text-to-video synthesis models demonstrate coherent, photorealistic generation of complex videos from a text description. However, most existing models lack fine-grained control over camera movement, which is critical for downstream applications related to content creation, visual effects, and 3D vision. Recently, new methods demonstrate the ability to generate videos with controllable camera poses—these techniques leverage pre-trained U-Net-based diffusion models that explicitly disentangle spatial and temporal generation. Still, no existing approach enables camera control for new, transformer-based video diffusion models that process spatial and temporal information jointly. Here, we propose to tame video transformers for 3D camera control using a ControlNet-like conditioning mechanism that incorporates spatiotemporal camera embeddings based on Plucker coordinates. The approach demonstrates state-of-the-art performance for controllable video generation after fine-tuning on the RealEstate10K dataset. To the best of our knowledge, our work is the first to enable camera control for transformer-based video diffusion models.
Figure 1: 3D camera control for text-to-video generation. We introduce a method that can control camera poses for text-to-video generation using video diffusion transformers. (left) The method takes as input a set of camera poses used to generate each frame of a rendered video. (center, right) Applying multiple camera trajectories with the same text prompt enables synthesis of complex scenes from varied set of viewpoints.
@@ 1 Introduction
Figure 2: Comparing text-to-video, text-to-4D, and camera-conditioned text-to-video generation. Text-to-4D approaches, such as 4D-fy [4] (top) have full control over camera through use of a 3D representation, but they lack photorealism. (middle) Methods for text-to-video generation [72] create realistic videos, but do not provide explicit control over the viewpoint. In contrast, camera-conditioned text-to-video generation (bottom) bridges the gap between the two paradigms by extending text-to-video generators with 3D camera control without using an explicit 3D representation. Please see the supplementary webpage for the corresponding video results.
@@
Text-to-video foundation models achieve unprecedented visual quality [8, 90]. They are trained on massive collections of images and videos and learn to synthesize remarkably consistent and physically plausible visualizations of the world. Yet, they lack built-in mechanisms for explicit 3D control during the synthesis process, requiring users to manipulate outputs through prompt engineering and trial and error—a slow, laborious, and computationally expensive process. For example, as Fig. 2 shows, state-of-the-art video models struggle to follow even simple “zoom-in” or “zoom-out” camera trajectories using text prompt instructions (see supplemental webpage). This lack of controllability limits interactivity and makes existing video generation techniques challenging to use for artists or other end users. We augment 2D video generation models with control over the position and orientation of the camera, providing finer-grained control compared to text prompting, and facilitating use of video generation models for downstream applications.
@@
Several contemporary works [119, 29, 135] propose methods for camera control of state-of-the-art, open-source video diffusion models. The key technical insight proposed by these methods is to add camera control by fine-tuning the temporal conditioning layers of a U-Net-based video generation model on a dataset with high-quality camera annotations. While these techniques achieve promising results, they are not applicable to more recent, high-quality transformer-based architectures [109, 78], such as Sora [8], SnapVideo [72], and Lumina-T2X [22], as these latest works simply do not have standalone temporal layers amenable to camera conditioning.
@@
Large video transformers represent a video as a (possibly compressed) sequence of tokens, applying self-attention layers to all the tokens jointly [8, 72]. Consequently, they do not have standalone temporal layers, which are essential for current camera conditioning approaches to integrate viewpoint information, making them inapplicable to such models. As the community shifts towards large video transformers to jointly model spatiotemporal dependencies in the data, it is critical to develop methods that provide similar capabilities for camera control. Our work designs a camera conditioning method tailored to the joint spatiotemporal computation used in large video transformers and takes a step towards taming them for controllable video synthesis.
@@
We develop our work on top of our implementation of SnapVideo [72], a state-of-the-art video diffusion model, that employs FIT-blocks [16] for efficient video modeling in the compressed latent space. We investigate various camera conditioning mechanisms in the fine-tuning scenario and explore trade-offs in terms of visual quality preservation and controllability. Our findings reveal that simply adapting existing approaches to video transformers does not yield satisfactory results: they either enable some limited amount of control while reducing the visual quality of the output video, or they entirely fail to control camera motion. Our key technical insight is to enable the control through spatiotemporal camera embeddings, which we derive by combining Plucker coordinates with the network input through a separately trained cross-attention layer. To the best of our knowledge, our work is the first to explore a ControlNet-like [150] conditioning mechanism for spatiotemporal transformers.
@@
We evaluate the method on a collection of manually crafted text prompts and unseen camera trajectories and compare to baseline approaches that incorporate previous camera control methods into a video transformer. Our approach achieves state-of-the-art results in terms of camera controllability and video quality, and also enables downstream applications such as multi-view, text-to-video generation, as depicted in Fig. 1. In contrast to existing image-to-3D methods (e.g., [65, 82, 110]), which are limited to object-centric scenes, our approach synthesizes novel views for real input images with complex environments.
@@
  • We propose a new method to tame large video transformers for 3D camera control. Our approach uses a ControlNet-like conditioning mechanism that incorporates spatiotemporal camera embeddings based on Plucker coordinates.
@@
  • We demonstrate state-of-the-art results in controllable video generation by applying the proposed conditioning method and fine-tuning scheme to the SnapVideo model [72].
@@ 2 Related work
Our method is connected to techniques related to text-to-video, text-to-3D, and text-to-4D generation. As this is a popular and fast-moving field, this section provides only a partial overview with a focus on the most relevant techniques; we refer readers to Po et al. [79] and Yunus et al. [145] for a more thorough review of related techniques.
@@
Text-to-video generation. Our work builds on recent developments in 2D video generation models. One such class of these techniques works by augmenting text-to-image models with layers that operate on the temporal dimension to facilitate video generation [7, 93, 123, 27, 6]. Video generation methods can be trained in a hybrid fashion on both images and video to improve the generation quality [5, 115, 133, 33, 27, 31, 116, 155]. While these techniques are primarily based on convolutional, U-Net-style architectures, a recent shift towards transformer-based models enables synthesis of much longer videos with significantly higher quality [8, 71, 72]. Still, these recent techniques do not enable synthesis with controllable camera motion.
@@
4D generation. Previous methods also tackle the problem of 4D generation, i.e., generating videos of dynamic 3D scenes from controllable viewpoints, usually from an input text prompt or image. Since the initial work on this topic using large-scale generative models [94], significant improvements in the visual quality and motion quality of generated scenes have been achieved [86, 63, 4, 154, 23, 134, 45, 1, 151, 127, 73, 55]. While these methods generate scenes based on text conditioning, other approaches convert an input image or video to a dynamic 3D scene [86, 153, 139, 77, 154, 63, 23, 146, 17, 124, 136, 117, 21, 97, 148, 142, 87, 54, 57, 108, 107, 40, 9, 59, 149]. Another line of work [2, 132] extends 3D GANs into 4D by training on 2D videos, however the quality is limited and models are trained on single category datasets. Still, all of these methods are focused on object-centric generation, typically based on 3D volumetric representations. As such, they typically do not incorporate background elements, and overall, they do not approach the level of photorealism demonstrated by the state-of-the-art video generation models used in our technique.
@@
Controllable generation with diffusion models. Methods for controllable generation using diffusion models have had significant impact, both in the context of image and video generation. For example, existing techniques allow controllable image generation conditioned on text, depth maps, edges, pose, or other signals [150, 137]. Furthermore, there is a line of work for 3D generation which conditions diffusion models on camera poses for view-consistent multi-view generation [120, 105, 11, 143, 52, 75, 24]. Our approach is most similar to related techniques in video generation that seek to control the camera position. For example, MotionCtrl [119] designs camera and object control mechanisms for the VideoCrafter1 [13] and SVD [6] models.
@@
Concurrent 3D camera control methods. Concurrent approaches enable camera control by conditioning the temporal layers of the network with camera pose information, e.g., using Plucker coordinates [29, 27, 128, 51] or other embeddings [135]. Interestingly, it is also possible to incorporate camera control into video generation models without additional training through manipulation and masking of attention layers, though this requires additional tracking, segmentation, or depth for each input video [39, 125, 38]. Another recent work [64] transfers motion, including camera motion, to other generated videos.
@@
Although these approaches demonstrate promising results for U-Net-based video diffusion models, the techniques are not applicable to modern video transformers that model spatio-temporal dynamics jointly. While another concurrent work [121] uses a transformer-based architecture for space and time, it does not tackle text-based generation for dynamic scenes but mainly novel view synthesis for 3D scenes. In our work, we design an efficient mechanism that enables camera control in video diffusion transformers using a ControlNet inspired mechanism without sacrificing visual quality.
@@ 3 Method
Figure 3: Overview of architecture. We adapt the Snap Video FIT architecture [72] to incorporate camera control. We take as input the noisy input video $\tilde{\bm{x}}$, camera extrinsics $\bm{C}_{f}$, and camera intrinsics $\bm{K}_{f}$ for each video frame $f$. Using the camera parameters, we compute the Plücker coordinates for each pixel within the video frames. Both the input video and Plücker coordinate frames are converted to patch tokens, and we condition the video patch tokens using a mechanism similar to ControlNet [150] (“Plücker conditioning” block). Then, the model estimates the denoised video $\hat{\bm{x}}$ by recurrent application of FIT blocks [16]. Each block reads information from the patch tokens into a small set of latent tokens on which computation is performed. The results are written to the patch tokens in an iterative denoising diffusion process.
@@ 3.1 Large text-to-video transformers
Text-to-video generation. Diffusion models have emerged as the dominant paradigm for large-scale video generation [34, 33, 8]. The standard setup considers the conditional distribution $p(\bm{x}|\bm{y})$ of videos $\bm{x}\in\mathbb{R}^{F\times H\times W}$ (consisting of $F$ frames of $H\times W$ resolution) given their textual descriptions $\bm{y}\in\mathcal{Y}^{L}$, consisting of $L$ (possibly padded) tokens from the alphabet $\mathcal{Y}$. Following [47], our video diffusion framework assumes a denoising model $D_{\bm{\theta}}:(\tilde{\bm{x}};\bm{y},\sigma)\mapsto\hat{\bm{x}}$ that predicts a clean video $\hat{\bm{x}}$ from the corresponding noised input $\tilde{\bm{x}}=\bm{x}+\sigma\bm{\varepsilon}$, $\sigma\in\mathbb{R}_{+},\bm{\varepsilon}\sim\mathcal{N}(\bm{0},\bm{I})$. The model is parametrized by a neural network $F_{\bm{\theta}}(\tilde{\bm{x}};\bm{y},\sigma)$ as $D_{\bm{\theta}}(\tilde{\bm{x}};\bm{y},\sigma)=c_{\text{out}}(\sigma)F_{\bm{\theta}}(c_{\text{in}}(\sigma)\tilde{\bm{x}};\bm{y},\sigma)+c_{\text{skip}}(\sigma)\tilde{\bm{x}}$. The optimization objective is defined as follows:
@@
$\mathcal{L}(\bm{\theta})\triangleq\underset{(\bm{x},\bm{y}),\sigma,\bm{\varepsilon}}{\mathbb{E}}\left[\|D_{\bm{\theta}}(\bm{x}+\sigma\bm{\varepsilon};\bm{y},\sigma)-\bm{x}\|_{2}^{2}\right]\longrightarrow\min_{\bm{\theta}}$ (1)
@@
We refer a reader to [72] and [47] for further details on the diffusion setup, which we adopted without modifications.
@@
Spatiotemporal transformers. Following SnapVideo [72], our video generator consists of two models: the base 4B-parameters generator, operating on 16-frames $36\times 64$ resolution videos, and a $288\times 512$ upsampler. The latter, a diffusion model itself, is fine-tuned from the base model and conditioned on the low-resolution videos. Each model uses FIT transformer blocks [16, 41] for efficient self-attention operations (see Fig. 3). An FIT model consists of $B$ blocks (we have $B=6$ in all the experiments) and first partitions each frame in an input video into patches [19] of resolution $h_{p}\times w_{p}$ (we use $h_{p}=w_{p}=4$ in all the experiments). These video patches are then independently projected via a feedforward (FF) layer to obtain a sequence of video tokens $[\bm{v}_{\ell}]_{\ell=1}^{L}\triangleq(\bm{v}_{1},...,\bm{v}_{L})\in\mathbb{R}^{L\times d}$ of length $L=F\times(H/h_{p})\times(W/w_{p})$ and dimensionality $d$. Next, each FIT block “reads” the information from this video sequence into a much shorter sequence of latent tokens $[\bm{z}_{m}]_{m=1}^{M}\triangleq(\bm{z}_{1},...,\bm{z}_{L})$ through a “read” cross-attention layer, followed by a feedforward layer. The core processing with self-attention layers is performed in this latent space, and then the result is written back to the video tokens through a corresponding “write” cross-attention layer (also followed by an FF layer). The latent tokens in each next FIT block are initialized from the previous one, which helps to propagate the computational results throughout the network. In this way, the entire computation occurs jointly in both spatial and temporal axes, which yields superior scalability properties [72]. However, it abandons the decomposed spatial/temporal computation nature of modern video diffusion U-Nets, which is vital for modern camera conditioning techniques [119, 135] to enforce control without compromising visual quality.

### v3

@@ Abstract
Modern text-to-video synthesis models demonstrate coherent, photorealistic generation of complex videos from a text description. However, most existing models lack fine-grained control over camera movement, which is critical for downstream applications related to content creation, visual effects, and 3D vision. Recently, new methods demonstrate the ability to generate videos with controllable camera poses—these techniques leverage pre-trained U-Net-based diffusion models that explicitly disentangle spatial and temporal generation. Still, no existing approach enables camera control for new, transformer-based video diffusion models that process spatial and temporal information jointly. Here, we propose to tame video transformers for 3D camera control using a ControlNet-like conditioning mechanism that incorporates spatiotemporal camera embeddings based on Plücker coordinates. The approach demonstrates state-of-the-art performance for controllable video generation after fine-tuning on the RealEstate10K dataset. To the best of our knowledge, our work is the first to enable camera control for transformer-based video diffusion models.
Figure 1: 3D camera control for text-to-video generation. We introduce a method that can control camera poses for text-to-video generation using video diffusion transformers. (left) The method takes as input a set of camera poses used to generate each frame of a rendered video. (center, right) Applying multiple camera trajectories with the same text prompt enables synthesis of complex scenes from a varied set of viewpoints. Video results: https://snap-research.github.io/vd3d.
@@ 1 Introduction
Figure 2: Comparing text-to-video, text-to-4D, and camera-conditioned text-to-video generation. We show time progressing across columns for generated videos from different approaches. Text-to-4D approaches, such as 4D-fy (Bahmani et al., 2024b) (top), have complete control over the camera through a 3D representation but lack photorealism. We visualize time progressing while simultaneously changing the viewpoint. (middle) Methods for text-to-video generation (Menapace et al., 2024) create realistic videos but do not provide explicit control over the viewpoint. In contrast, camera-conditioned text-to-video generation (bottom) bridges the gap between the two paradigms by extending text-to-video generators with 3D camera control without using an explicit 3D representation. Specifically, we condition VD3D with a zoom-in camera trajectory to control the generation.
@@
Text-to-video foundation models achieve unprecedented visual quality (Brooks et al., 2024; Sharma et al., 2024). They are trained on massive collections of images and videos and learn to synthesize remarkably consistent and physically plausible visualizations of the world. Yet, they lack built-in mechanisms for explicit 3D control during the synthesis process, requiring users to manipulate outputs through prompt engineering and trial and error—a slow, laborious, and computationally expensive process. For example, as Fig. 2 shows, state-of-the-art video models struggle to follow even simple “zoom-in” or “zoom-out” camera trajectories using text prompt instructions (see supplemental webpage). This lack of controllability limits interactivity and makes existing video generation techniques challenging to use for artists or other end users. We augment 2D video generation models with control over the position and orientation of the camera, providing finer-grained control compared to text prompting, and facilitating use of video generation models for downstream applications.
@@
Several contemporary works (Wang et al., 2023e; He et al., 2024a; Yang et al., 2024b) propose methods for camera control of state-of-the-art, open-source video diffusion models. The key technical insight proposed by these methods is to add camera control by fine-tuning the temporal conditioning layers of a U-Net-based video generation model on a dataset with high-quality camera annotations. While these techniques achieve promising results, they are not applicable to more recent, high-quality transformer-based architectures (Vaswani et al., 2017; Peebles & Xie, 2023), such as Sora (Brooks et al., 2024), SnapVideo (Menapace et al., 2024), and Lumina-T2X (Gao et al., 2024a), as these latest works simply do not have standalone temporal layers amenable to camera conditioning.
@@
Large video transformers represent a video as a (possibly compressed) sequence of tokens, applying self-attention layers to all the tokens jointly (Brooks et al., 2024; Menapace et al., 2024). Consequently, as diffusion transformers do not have standalone temporal layers, they are incompatible with current camera conditioning approaches. As the community shifts towards large video transformers to jointly model spatiotemporal dependencies in the data, it is critical to develop methods that provide similar capabilities for camera control. Our work designs a camera conditioning method tailored to the joint spatiotemporal computation used in large video transformers and takes a step towards taming them for controllable video synthesis.
@@
We develop our work on top of our implementation of SnapVideo (Menapace et al., 2024), a state-of-the-art video diffusion model, that employs Far-reaching Interleaved Transformers (FIT) blocks (Chen & Li, 2023) for efficient video modeling in the compressed latent space. We investigate various camera conditioning mechanisms in the fine-tuning scenario and explore trade-offs in terms of visual quality preservation and controllability. Our findings reveal that simply adapting existing approaches to video transformers does not yield satisfactory results: they either enable some limited amount of control while reducing the visual quality of the output video, or they entirely fail to control camera motion. Our key technical insight is to enable the control through spatiotemporal camera embeddings, which we derive by combining Plücker coordinates with the network input through a separately trained cross-attention layer. To the best of our knowledge, our work is the first to explore a ControlNet-like (Zhang et al., 2023) conditioning mechanism for spatiotemporal transformers.
@@
We evaluate the method on a collection of manually crafted text prompts and unseen camera trajectories and compare to baseline approaches that incorporate previous camera control methods into a video transformer. Our approach achieves state-of-the-art results in terms of camera controllability and video quality, and also enables downstream applications such as multi-view, text-to-video generation, as depicted in Figure 1. In contrast to existing image-to-3D methods (e.g., (Liu et al., 2024a; Qian et al., 2024b; Voleti et al., 2024a)), which are limited to object-centric scenes, our approach synthesizes novel views for real input images with complex environments.
@@
  • We propose a new method to tame large video transformers for 3D camera control. Our approach uses a ControlNet-like conditioning mechanism that incorporates spatiotemporal camera embeddings based on Plücker coordinates.
@@
  • We show state-of-the-art results in camera-controllable video synthesis by applying the proposed conditioning method and fine-tuning scheme to the SnapVideo-based model (Menapace et al., 2024).
@@ 2 Related work
Our method is connected to techniques related to text-to-video, text-to-3D, and text-to-4D generation. As this is a popular and fast-moving field, this section provides only a partial overview with a focus on the most relevant techniques; we refer readers to  Po et al. (2023) and  Yunus et al. (2024) for a more thorough review of related techniques.
@@
Text-to-video generation. Our work builds on recent developments in 2D video generation models. One such class of these techniques works by augmenting text-to-image models with layers that operate on the temporal dimension to facilitate video generation (Blattmann et al., 2023b; Singer et al., 2023a; Wu et al., 2023; Guo et al., 2024; Blattmann et al., 2023a). Video generation models can be trained in a hybrid fashion on both images and videos to improve the generation quality (Bain et al., 2021; Wang et al., 2023b; Xue et al., 2022; Ho et al., 2022a; Guo et al., 2024; He et al., 2022; Wang et al., 2023c; Zhou et al., 2022). While they are primarily based on convolutional, U-Net-style architectures, a recent shift towards transformer-based architectures enables synthesis of much longer videos with significantly higher quality (Brooks et al., 2024; Ma et al., 2024b; Menapace et al., 2024; Ma et al., 2024a). Still, these methods do not enable synthesis with controllable camera motion.
@@
4D generation. Previous methods also tackle the problem of 4D generation, i.e., generating videos of dynamic 3D scenes from controllable viewpoints, usually from an input text prompt or image. Since the initial work on this topic using large-scale generative models (Singer et al., 2023b), significant improvements in the visual quality and motion quality of generated scenes have been achieved (Ren et al., 2023; Ling et al., 2024a; Bahmani et al., 2024b; Zheng et al., 2024a; Bahmani et al., 2024a). While these methods generate scenes based on text conditioning, other approaches convert an input image or video to a dynamic 3D scene (Ren et al., 2023; Zhao et al., 2023; Yin et al., 2023; Pan et al., 2024; Zheng et al., 2024a; Ling et al., 2024a; Gao et al., 2024b; Zeng et al., 2024; Chu et al., 2024). Another line of work (Bahmani et al., 2023a; Xu et al., 2023) extends 3D GANs into 4D by training on 2D videos, however the quality is limited and models are trained on single category datasets. All of these methods are focused on object-centric generation, typically based on 3D volumetric representations. As such, they typically do not incorporate background elements, and overall, they do not approach the level of photorealism demonstrated by the state-of-the-art video generation models used in our technique (see Fig. 2).
@@
Controllable generation with diffusion models. Methods for controllable generation using diffusion models have had significant impact, both in the context of image and video generation. For example, existing techniques allow controllable image generation conditioned on text, depth maps, edges, pose, or other signals (Zhang et al., 2023; Ye et al., 2023). While Chen et al. (2024) developed a ControlNet-based mechanism for transformer-based diffusion, limited to spatial conditioning with spatial signals, we explore conditioning mechanisms for camera poses (a spatio-temporal signal with an intricate temporal component) in a spatio-temporal transformer. Furthermore, there is a line of work for 3D generation that conditions diffusion models on camera poses for view-consistent multi-view generation (Watson et al., 2023; Tseng et al., 2023; Chan et al., 2023; Yu et al., 2023a; Kumari et al., 2024; Müller et al., 2024; Gao et al., 2024c). Our approach is most similar to related techniques in video generation that seek to control the camera position. For example, MotionCtrl (Wang et al., 2023e) designs camera and object control mechanisms for the VideoCrafter1 (Chen et al., 2023b) and SVD (Blattmann et al., 2023a) models. However, MotionCtrl is designed for U-Net-based approaches and does not directly apply to video diffusion transformers.
@@
Concurrent 3D camera control methods. Concurrent approaches enable camera control by conditioning the temporal layers of the network with camera pose information, e.g., using Plücker coordinates (He et al., 2024a; Guo et al., 2024; Xu et al., 2024b; Kuang et al., 2024) or other embeddings (Yang et al., 2024b). Interestingly, it is also possible to incorporate camera control into video generation models without additional training through manipulation and masking of attention layers, though this requires additional tracking, segmentation, or depth for each input video (Hu et al., 2024; Xiao et al., 2024; Hou et al., 2024). Another recent work (Ling et al., 2024b) transfers motion, including camera motion, to other generated videos.
@@
Although these approaches demonstrate promising results for U-Net-based video diffusion models, the techniques are not applicable to modern video transformers that model spatio-temporal dynamics jointly. While another concurrent work (Watson et al., 2024) uses a transformer-based architecture for space and time, it does not tackle text-based generation for dynamic scenes but focuses on novel view synthesis from an input image. In our work, we design an efficient mechanism that enables camera control in video diffusion transformers using a ControlNet inspired mechanism without sacrificing visual quality.
@@ 3 Method
Figure 3: Overview of architecture. We adapt a FIT-based architecture (Chen & Li, 2023) to incorporate camera control. We take as input the noisy input video $\tilde{\bm{x}}$, camera extrinsics $\bm{C}_{f}$, and camera intrinsics $\bm{K}_{f}$ for each video frame $f$. We compute the Plücker coordinates for each pixel within the video frames using the camera parameters. Both the input video and Plücker coordinate frames are converted to patch tokens, and we condition the video patch tokens using a mechanism similar to ControlNet (Zhang et al., 2023) (“Plücker conditioning” block). Then, the model estimates the denoised video $\hat{\bm{x}}$ by recurrent application of FIT blocks (Chen & Li, 2023). Each block reads information from the patch tokens into a small set of latent tokens on which computation is performed. The results are written to the patch tokens in an iterative denoising diffusion process.
@@ 3.1 Large text-to-video transformers
Text-to-video generation. Diffusion models have emerged as the dominant paradigm for large-scale video generation (Ho et al., 2022b; a; Brooks et al., 2024). The standard setup considers the conditional distribution $p(\bm{x}|\bm{y})$ of videos $\bm{x}\in\mathbb{R}^{F\times H\times W}$ (consisting of $F$ frames of $H\times W$ resolution) given their text descriptions $\bm{y}\in\mathcal{Y}^{L}$, consisting of $L$ (possibly padded) tokens from the alphabet $\mathcal{Y}$. Following (Karras et al., 2022), our video diffusion framework assumes a denoising model $D_{\bm{\theta}}:(\tilde{\bm{x}};\bm{y},\sigma)\mapsto\hat{\bm{x}}$ that predicts a clean video $\hat{\bm{x}}$ from the corresponding noised input $\tilde{\bm{x}}=\bm{x}+\sigma\bm{\varepsilon}$, where $\bm{\varepsilon}\sim\mathcal{N}(\bm{0},\bm{I})$ is standard Gaussian noise and $\sigma\sim\log\mathcal{N}(P_{\text{mean}},P_{\text{std}}^{2})$ is the noise strength, sampled from log-normal distribution with location $P_{\text{mean}}$ and scale $P_{\text{std}}$. The model is parametrized by a neural network $F_{\bm{\theta}}(\tilde{\bm{x}};\bm{y},\sigma)$ as $D_{\bm{\theta}}(\tilde{\bm{x}};\bm{y},\sigma)=c_{\text{out}}(\sigma)F_{\bm{\theta}}(c_{\text{in}}(\sigma)\tilde{\bm{x}};\bm{y},\sigma)+c_{\text{skip}}(\sigma)\tilde{\bm{x}}$, where $c_{\text{in}}(\sigma),c_{\text{out}}(\sigma)$ and $c_{\text{skip}}(\sigma)$ are input, output and residual scaling factors from (Karras et al., 2022). The minimization objective is defined as
@@
$\mathcal{L}(\bm{\theta})=\underset{p(\bm{x},\bm{y},\sigma,\bm{\varepsilon})}{\mathbb{E}}\left[\|D_{\bm{\theta}}(\bm{x}+\sigma\bm{\varepsilon};\bm{y},\sigma)-\bm{x}\|_{2}^{2}\right].$ (1)
@@
We refer the reader to (Menapace et al., 2024) and (Karras et al., 2022) for further details on the diffusion setup, which we adopted without modifications.
@@
Spatiotemporal transformers. Following SnapVideo (Menapace et al., 2024), our video generator consists of two models: the base 4B-parameters generator, operating on 16-frames $36\times 64$ resolution videos, and a $288\times 512$ upsampler. The latter, a diffusion model itself, is fine-tuned from the base model and conditioned on the generated low-resolution videos. Each model uses FIT transformer blocks (Chen & Li, 2023; Jabri et al., 2022) for efficient self-attention operations (see Fig. 3). An FIT model consists of $B$ blocks (we have $B=6$ in all the experiments) and first partitions each frame in an input video into patches (Dosovitskiy et al., 2021) of resolution $h_{p}\times w_{p}$ (we use $h_{p}=w_{p}=4$ in all the experiments). These video patches are then independently projected via a feedforward (FF) layer to obtain a sequence of video tokens $\bm{v}^{(b)}_{\ell:L}\triangleq(\bm{v}_{1},...,\bm{v}_{L})\in\mathbb{R}^{L\times d}$ of length $L=F\times(H/h_{p})\times(W/w_{p})$ and dimensionality $d$. Next, each FIT block “reads” the information from this video sequence into a much shorter sequence of $M$ latent tokens $\bm{z}^{(b)}_{m:M}\triangleq(\bm{z}_{1},...,\bm{z}_{M})$ through a “read” cross-attention layer, followed by a feedforward layer. The core processing with self-attention layers is performed in this latent space, and then the result is written back to the video tokens through a corresponding “write” cross-attention layer (also followed by an FF layer). The latent tokens in each subsequent FIT block are initialized from the previous one, which helps to propagate the computational results throughout the network. In this way, the entire computation occurs jointly in both spatial and temporal axes, which yields superior scalability (Menapace et al., 2024). However, it abandons the decomposed spatial/temporal computation nature of modern video diffusion U-Nets, which modern camera conditioning techniques (Wang et al., 2023e; Yang et al., 2024b) rely on to enforce control without compromising visual quality.